OpenStudy (anonymous):

what is the derivative of 3x^-1/2

OpenStudy (anonymous):

\[-6x ^{(-\frac{ 3 }{ 2 })}\]

OpenStudy (anonymous):

@Mr.Me can u look over my answer and see if its correct

OpenStudy (anonymous):

heres the problem, \[\sqrt{x} +3/ \sqrt{x}\]

OpenStudy (anonymous):

after finding the derivative i got \[\frac{ 1 }{ 2\sqrt{x} } + -6x ^{-3/2}\]

OpenStudy (anonymous):

yeah good

OpenStudy (anonymous):

its-3/2 in the second one right?

OpenStudy (anonymous):

should be -6x^(-3/2)

OpenStudy (anonymous):

yea \[\frac{ 3 }{ \sqrt{x} }\]

OpenStudy (anonymous):

i wont distribute the minus sign in the answer right?

OpenStudy (anonymous):

so is my answer correct?

zepdrix (zepdrix):

 I think your 2 is in the wrong spot :D